A 0.50 Kg ball with a speed of 12m hits a wvall and only 0.75 of the original balls collision with the ball? PROBLEM!2i speed. . How much kinetic energy is lost in the 8 Points a: What is the initial momentum of the ball and its direction? b What is the final momentum of the ball and its direction? C: What is the initial kinetic energy of the ball? d: How much kinetic energy has been lost in the balls collision with the ball?
Answer #1

Final velocity will be,

v=12*0.75

v=9m/s

a)

Initial momentum is mu

=0.5*12

=6 kg-m/s

b)

Final momentum will be

=0.5*9

=4.5 kg-m/s

c)

Initial kinetic energy is

=0.5*0.5*12*12

36J

d) final kinetic energy is

0.5*0.5*9*9

=20.25J

So energy lost is

36-20.25

=15.75J